Read Frosty Drew Observatory and Science Center's Update on SARS-CoV-2 / Coronavirus Disease 2019 and our Reopening Plan. Updated: August 5, 2020

Log, Feb 4, 2000

2 people. Joe went to the dome to put the repaired cable on the scope. The good news is that we've fixed the runaway dec motor problem. The bad news is that the movement in that axis is decidedly spotty. More work on that mechanism is required to get action smoothed out.

We actually had a visitor tonight! An 8" LX-200 owner from New Bedford came by after hearing about us a week or two back. He had to see the 16"! Unfortunately, that was all he could see as there was complete cloud cover followed by snow flurries.

-Les Coleman

Leslie Coleman
Author:
Leslie Coleman
Entry Date:
Feb 4, 2000
Published Under:
Leslie Coleman's Log
Subscribe to Leslie Coleman's Log RSS Feed